a ball is projected horizontally from the top of a vertical building 25m above the ground level with an initial velocity of 8.25m/s.

a. how long is the ball in flight before hiiting the ground

b. how far from the building does the ball strike the ground


1
Expert's answer
2021-11-08T17:03:38-0500

a) "h=gt^2/2\ \to t=\sqrt{2h/g}=\sqrt{2\cdot25/9.8}=2.56\ (s)"


b) "l=vt=8.25\cdot2.56=21.12\ (s)"